# Understanding intra-node communication in HPC systems and Datacenters

### v1

@@ 3.1 Baseline end-node architecture
The CELLIA cluster comprises several computing nodes with GPUs and NVMe disks. Specifically, each one of these end nodes has two CPUs, Intel Xeon Silver 4116, connected to the PCIe Gen3, and $192$GB of RAM. These nodes also have an EDR InfiniBand network interface (a.k.a. HCA) MCX556A-ECAT ConnectX-5 dual-port. The InfiniBand EDR technology provides a data rate of 100 $\mathrm{G}\mathrm{b}\mathrm{p}\mathrm{s}$ per HCA port (i.e., $12.5$ GB/s), with packets of $4$ $\mathrm{KiB}$ of MTU. These packets contain a $64$B header and a $4032$B payload. This cluster has a management 1GbE network, so each node has a network interface (NIC) for management purposes. In addition, each node has installed an NVIDIA Tesla T4 GPU over PCIe Gen3 x16 and an NVMe disk Ultrastar SN200 Series NVMe SSD over PCIe Gen3 x8. This versatility allows the node to perform a wide range of tasks, and we aim to model each component comprehensively to simulate various scenarios effectively.
Note that this heterogeneous node architecture has several devices attached to the PCIe (e.g., the CPU, GPU, NVMe, and HCA), so the intra-node network could be a bottleneck when inter-node communication involves other devices placed in different nodes. In the following section, we describe the characterization of the communication involving the intra-node network based on PCIe and the inter-node network.
@@ 3.2 PCIe communication characterization
To accurately model the PCIe subsystem communication, we first need to understand the specific characteristics of the PCI-Express system. We assume PCI-Express version 3.0, which can transfer data at 8 Gbps per lane. Since our HCA [^1] We use the term Host Channel Adapter (HCA) from the InfiniBand jargon to distinguish it from the 1GbE NIC in our assumed end-node model. transmits over 16 PCI lanes, the total data transfer rate from the intra-node network to the HCA can reach up to 128 Gbps (i.e., $16$GB/s). Considering the PCIe version and its $128$b/$130$b encoding, we can transfer accurate data close to $126$ Gbps. At the same time, the HCA can inject traffic into the inter-node network at a maximum rate of $100$Gbps since it is InfiniBand EDR technology. Thus, the intra-node communication may become a bottleneck when the HCA is saturated. Figure 2 shows the number of PCIe lanes and maximum payload size (MPS) for the different end-node devices. We assume that our PCIe communication model adjusts the MPS to $128$B.
@@
We assume that all end-node devices are connected through a PCIe Root Complex (RC) without using PCIe switches (see Figure 1). This PCIe configuration can impact communication performance between two different endpoints (E1 and E2), as the communication model follows the following steps: $EP1\rightarrow RC\rightarrow CPU\rightarrow RC\rightarrow EP2$. As end nodes in our real cluster follow this approach, and we want to validate our simulation model compared to these nodes, we have modeled this PCIe communication. Note that the CPU and RC involvement in the communication impact performance, especially when communicating between two end-node devices that do not require CPU involvement. The most efficient path typically uses a PCIe Switch, through which these devices communicate directly. By contrast, the CPU can communicate with the HCA in our tests by traversing the RC only once.
@@
$\quad NumberACKs\times DLPPTime$
@@ 3.3 Generic intra-node network model
Apart from incorporating the PCIe model described before, we have extended the SAURON network simulator [15]. to model a generic intra-node network and realistic communication patterns. SIM is a highly scalable simulator that models various inter-node communication patterns, topologies, and intra- and inter-node network configurations. Figure 3 shows the end-node generic architecture modeled in the SIM simulator.
Figure 3: SAURON simulator host architecture.
We assume a configurable number of accelerators and other functional units (e.g., CPUs or storage devices) interconnected “all-to-all” using a switch in the intra-node interconnection network. This switch can be configured to provide the desired bandwidth so the intra-node network can process the incoming traffic addressed to the end-node devices, and the outgoing traffic they generate. Indeed, we can also configure the bandwidth between this switch and the end-node NIC, i.e., the bandwidth between the intra- and inter-node network. Note that this generic model permits configuring the intra-node network according to emerging technologies (e.g., PCIe Gen4, Gen5, Gen6, NVLink, Habana Gaudi, etc.). Indeed, we can configure the number of lanes, the bandwidth for the different end-node devices, their maximum payload size (MPS), etc.
We assume messages are generated at accelerators (GPUs, TPUs, etc.) and functional units (CPUs, NVMe, etc.). Messages at these devices are split into packets (according to the intra-node network MPS) and injected into the intra-node network. Note that these packets can be addressed to other devices in the same or different end nodes. In the latter case, NICs convert the intra-node packets to inter-node packets and inject them into the inter-node network. This conversion involves populating inter-node packets, with a size (i.e., MTU) different from that of intra-node packets, with the header and payloads of inter-node network packets. Therefore, this conversion may introduce additional latency.
As described in the following section, we can configure the message size at accelerators and functional units, the destination distribution, and the generation rate in our model. This flexibility permits reproducing the communication patterns observed in the training phase of LLMs when several types of parallelism are used (e.g., DP, TP, or PP).
@@ 3.4 Generic LLM communication modeling
We have modeled network traffic generated by LLMs, specifically during the training phase. To accurately capture LLM communication traffic characteristics, we have considered different approaches. First, AI training follows a cyclic pattern, where communication operations repeat across data batches provided to the LLM. In this context, it is essential to distinguish between Data Parallelism (DP) and Model Parallelism (MP), as described in Section 2.4. Between these communication operations, accelerators (e.g., GPUs) perform computation tasks until reaching different sub-layers (e.g., Multi-Head Attention (MHA), Feed-forward Neural Network (FFN)), which then trigger new communication operations. In our case, all accelerators have identical computational capabilities, processing the same amount of data at the same speed. As a result, all accelerators initiate their next communication operation simultaneously.
@@
We assume the LLM is already stored in the accelerators’ memory, eliminating the need to exchange messages with external memory resources (e.g., NVMe disks). Instead, all communication occurs between accelerators to share information during the Forward and Backward passes in MP or to update gradients in DP. Our LLM communication model defines five distinct traffic patterns: $C1$, $C2$, $C3$, $C4$, and $C5$. From $C1$ to $C4$, we model a massive LLM that requires MP to distribute the model across multiple accelerators. The key difference between these traffic patterns is the predominant form of MP: Tensor Parallelism (TP) or Pipeline Parallelism (PP).
@@
$C1$ assumes that TP is used extensively, i.e., most sub-layer communications follow a TP pattern (e.g., AllReduce, AllGather operations). This scenario generates significant traffic due to the need to update gradients during forward and backward propagation between devices (see Section 2.4). As a result, this traffic pattern generates $20\%$ of the communication to the inter-node network and $80\%$ to the intra-node network. Next, $C2$ and $C3$ also employ MP; however, the transferred data size between devices is smaller as these configurations rely more on PP than TP. This reduces inter-node traffic, as PP primarily uses Point-to-Point operations. Specifically, these traffic patterns generate $15\%$ and $10\%$ of the traffic to the inter-node network, while $85\%$ and $90\%$ remain within the intra-node network, respectively. Finally, $C4$ uses MP without TP, relying exclusively on PP. In this case, inter-node communication is minimized (e.g., $5\%$), as fewer updates are required between stages. Consequently, intra-node network communication exceeds $95\%$.
@@
By contrast, $C5$ does not implement MP. In this scenario, we assume the model is not large enough to require multiple accelerators. Since the model fits within a single accelerator, only DP is used to accelerate training, and no inter-node communication is needed. As a result, all communication occurs within the same node ($100\%$ intra-node communication).
@@ 4 Evaluation
This section describes the experiments conducted to validate our simulation model compared to a real cluster environment. We also show the results of scale-out simulations using realistic intra- and inter-node network configurations and communication patterns.
@@ 4.1 Simulation model validation
This section describes the validation of the PCIe intra-node simulation model compared to the baseline end-node architecture (see Section 3.1). We have performed two different experiments using the OSU micro-benchmarks [16] and the Infiniband performance tests [17] (IB Perftests). For the OSU micro-benchmarks, we used a host-based point-to-point latency test (osu_latency) that measures the latency between two compute nodes during a ping-pong message exchange of a certain data size. Many iterations of this exchange are executed and the average one-way latency value is obtained. As blocking MPI_Send and MPI_Recv functions are used, the waiting overhead at the host side is included in the latency values. To reduce this overhead, we used the IB Perftests that generate direct communication for user-space applications directly accessing the InfiniBand hardware, so they do not introduce the overhead of specific communication libraries. More precisely, we performed ib_send_lat, ib_write_lat, and ib_read_lat tests using payloads ranging from 128 Bytes to 4 MiB. As the MPS of the end-node PCIe hardware is 128 Bytes, messages smaller than this size have identical latency when they fit within a single 128B-payload TLP, so we omitted these values. Table 1 shows the bandwidth results for these tests when varying the generated message size.
Table 1: Bandwidth results ($\mathrm{GiB}$/$\mathrm{s}$) when communicating two nodes in the real cluster.
Msg. Size osu_latency ib_read ib_write ib_send
128 $\mathrm{B}$ $0.54$ $0.37$ $0.44$ $0.41$
256 $\mathrm{B}$ $1.04$ $0.79$ $0.87$ $0.77$
512 $\mathrm{B}$ $2.04$ $1.51$ $1.75$ $1.64$
$\mathrm{KiB}$ $3.44$ $2.74$ $3.30$ $3.10$
$\mathrm{KiB}$ $6.17$ $6.63$ $7.35$ $6.22$
$\mathrm{KiB}$ $8.41$ $9.90$ $11.02$ $11.00$
$\mathrm{KiB}$ $10.39$ $11.38$ $11.58$ $11.55$
16 $\mathrm{KiB}$ $11.11$ $11.78$ $11.53$ $11.63$
32 $\mathrm{KiB}$ $11.64$ $11.80$ $11.60$ $11.67$
64 $\mathrm{KiB}$ $11.93$ $11.81$ $11.62$ $11.60$
128 $\mathrm{KiB}$ $12.08$ $12.09$ $11.90$ $11.90$
256 $\mathrm{KiB}$ $12.16$ $12.09$ $11.92$ $11.93$
512 $\mathrm{KiB}$ $12.20$ $12.09$ $11.93$ $11.92$
$\mathrm{MiB}$ $12.21$ $12.09$ $11.93$ $11.93$
$\mathrm{MiB}$ $12.17$ $12.06$ $11.93$ $11.94$
$\mathrm{MiB}$ $12.16$ $12.03$ $11.86$ $11.94$
As we can see, the maximum link bandwidth for the ib_read_lat (i.e., around $12.1$ out of $12.5$ GB/s) is first achieved with $128$KiB messages. Larger messages cannot reach a higher bandwidth due to the packet header overhead (i.e., $60$B), since the maximum payload for $4$KB MTU packets in the InfiniBand network is $4096-60=4036$. Anyway, the obtained link bandwidths are similar regardless whether OSU or IB Perftests are used.
Table 2 shows a similar trend in latency. OSU has higher latency than IB Perftest for intermediate message sizes, due to the blocking MPI_Send and MPI_Receive calls. As we have not included the host-side latency of specific communication libraries in the simulation model, we will focus hereafter on the IB Perftest tests.
@@
Msg. Size osu_latency ib_read ib_write ib_send
128 $\mathrm{B}$ $1.61$ $2.03$ $1.12$ $1.20$
256 $\mathrm{B}$ $2.09$ $2.07$ $1.56$ $1.59$
512 $\mathrm{B}$ $1.96$ $2.02$ $1.58$ $1.64$
$\mathrm{KiB}$ $2.20$ $2.15$ $1.70$ $1.77$
$\mathrm{KiB}$ $3.00$ $2.43$ $1.95$ $2.02$
$\mathrm{KiB}$ $3.90$ $2.88$ $2.46$ $2.56$
$\mathrm{KiB}$ $5.52$ $3.40$ $2.84$ $2.94$
16 $\mathrm{KiB}$ $7.42$ $4.28$ $3.88$ $3.86$
32 $\mathrm{KiB}$ $9.26$ $5.68$ $5.41$ $5.32$
64 $\mathrm{KiB}$ $14.14$ $8.38$ $8.06$ $7.97$
128 $\mathrm{KiB}$ $23.32$ $13.66$ $13.39$ $13.25$
256 $\mathrm{KiB}$ $26.41$ $24.25$ $24.27$ $24.10$
512 $\mathrm{KiB}$ $47.88$ $45.40$ $45.73$ $45.41$
$\mathrm{MiB}$ $91.85$ $87.73$ $88.95$ $88.46$
$\mathrm{MiB}$ $177.96$ $173.31$ $174.65$ $173.74$
$\mathrm{MiB}$ $350.68$ $343.93$ $345.97$ $344.31$
In sections 3.1 and 3.2, we described that the real cluster achieves speeds of over $12.5$ $\mathrm{GiB}$/$\mathrm{s}$ on the Infiniband inter-node network when transmitting raw data (i.e., without considering packet headers) and $16$ $\mathrm{GiB}$/$\mathrm{s}$ on the PCIe intra-node network. Therefore, the Infiniband network becomes the bottleneck when the intra-node network demands full speed to the inter-node network. Table 1 shows that maximum bandwidth is achieved with $128$KB messages (i.e., the PCIe MPS). For larger messages, latency skyrockets due to this bottleneck, as TLPs accumulate in the intra-node network while waiting to be split into InfiniBand packets.
@@
Figure 4: Experiment results in the real cluster and the simulator when communicating two end-node devices.
@@
We modeled the ib_write Infiniband test in our simulator to validate the PCIe intra-node model. Figure 4 shows the bandwidth and latency results obtained running the ib_write test both in the simulator and the real cluster. Figure 4(a) compares the bandwidth results, which are virtually identical between the two series. Figure 4(b) shows the latency results that follow the same trend. These results show that, in this simple scenario, the PCIe intra-node model essentially behaves as expected. It splits conveniently PCIe transactions into TLPs (i.e., intra-node packets) sent through the end-node NIC device, which compose a network packet each with MTU${}=4$KB. These packets are sent through the InfiniBand network and, when received at the destination NIC, they are transformed again into TLPs and forwarded to the receiving buffer at the host-side user space application, so the performance metrics are measured.

### v2

@@ 3.1 Baseline end node architecture
The CELLIA cluster comprises several computing nodes; each end node has two CPUs, Intel Xeon Silver 4116, connected to the PCIe Gen3, and 192192GB of RAM. These nodes also have an EDR InfiniBand network interface (a.k.a. HCA) MCX556A-ECAT ConnectX-5 dual-port. The InfiniBand EDR technology provides a data rate of 100 $\mathrm{G}\mathrm{b}\mathrm{p}\mathrm{s}$ per HCA port (i.e., $12.5$ GB/s), with packets of 44 $\mathrm{KiB}$ of MTU. These packets contain a 6464B header and a 40324032B payload. CELLIA has a management 1GbE network, so each node has a network interface (NIC) for management purposes. In addition, each node has installed an NVIDIA Tesla T4 GPU over PCIe Gen3 x16 and an NVMe disk Ultrastar SN200 Series NVMe SSD over PCIe Gen3 x8. Note that this heterogeneous node architecture has several devices attached to the PCIe (e.g., the CPU, GPU, NVMe, and HCA), so the intra-node network could be a bottleneck when inter-node communication involves other devices placed in different nodes.
@@ 3.2 PCIe communication characterization
To accurately model the PCIe subsystem communication, we first need to understand the specific characteristics of the PCI-Express system. We assume PCI-Express version 3.0, which can transfer data at 8 Gbps per lane. Since our HCA [^2] We use the term Host Channel Adapter (HCA) from the InfiniBand jargon to distinguish it from the 1GbE NIC in our assumed end node model. transmits over 16 PCI lanes, the total data transfer rate from the intra-node network to the HCA can reach up to 128 Gbps (i.e., 1616GB/s). Considering the PCIe version and its 128128b/130130b encoding, we can transfer accurate data close to $126$ Gbps. At the same time, the HCA can inject traffic into the inter-node network at a maximum rate of 100100Gbps since it is InfiniBand EDR technology. Thus, the intra-node communication may become a bottleneck when the HCA is saturated. Figure 2 shows the number of PCIe lanes and maximum payload size (MPS) for the different end-node devices. We assume that our PCIe communication model adjusts the MPS to 128128B.
@@
We assume that all end-node devices are connected through a PCIe Root Complex (RC) without using PCIe switches (see Figure 1). This PCIe configuration can impact communication performance between two different endpoints (E1 and E2), as the communication model follows the following steps: $E1\rightarrow RC\rightarrow CPU\rightarrow RC\rightarrow E2$. As end nodes in our real cluster follow this approach, and we want to validate our simulation model compared to these nodes, we have modeled this PCIe communication. Note that the CPU and RC involvement in the communication impact performance, especially when communicating between two end-node devices that do not require CPU involvement. The most efficient path typically uses a PCIe Switch, through which these devices communicate directly. By contrast, the CPU can communicate with the HCA in our tests by traversing the RC only once.
@@
$\qquad NumberACKs\times DLPPTime\$
@@ 3.3 Generic intra-node network model
Apart from integrating the PCIe model in our SAURON simulator [15], we have extended this tool to model a generic intra-node network and realistic communication patterns. SAURON is a highly scalable simulator modeling inter-node communication patterns, topologies, and routing configurations. Figure 3 shows the intra- and inter-node architecture modeled in the SAURON simulator.
Figure 3: Generic intra- and inter-node network architecture.
We assume a configurable number of accelerators and other functional units (e.g., CPUs or storage devices) interconnected “all-to-all” using a switch in the intra-node network. This switch can be configured to provide the desired bandwidth so the intra-node network can process the incoming traffic addressed to the end-node devices and the outgoing traffic they generate. Indeed, we can also configure the bandwidth between this switch and the end-node NIC, i.e., the bandwidth between the intra- and inter-node network. Note that this generic model permits configuring the intra-node network according to current and emerging technologies (e.g., PCIe 5.0, 6.0, 7.0, NVLink, UALink, etc.). Indeed, we can configure the number of lanes, the bandwidth for the different end-node devices, their maximum payload size (MPS), etc.
We assume messages are generated at accelerators (GPUs, TPUs, etc.) and functional units (CPUs, NVMe, etc.), which need to split them into packets (according to the intra-node network MPS) that are injected into the intra-node network. Intra-node packets can be addressed to other devices within the same end node or placed at different end nodes. In the latter case, end-node NICs convert the intra-node packets to inter-node packets and inject them into the inter-node network. This conversion involves populating inter-node packets with the headers and payloads of intra-node network packets, introducing additional latency and overhead.
As later described, in our simulation model, we can configure the message size at accelerators and functional units, the destination distribution, and the generation rate. This flexibility permits reproducing realistic communication patterns, such as those observed in the training phase of LLMs, where several types of parallelism are used.
@@ 3.4 Generic LLM communication modeling
In our simulator, we have included a generic network traffic model based on the communication operations generated by Large Language Models (LLMs), which are commonly used in Supercomputers and Data centers, as described in Section 2.4. Specifically, training in AI-based applications follows a cyclic pattern, where communication operations repeat across data batches provided to the LLM when splitting a dataset into chunks. During training, accelerators (e.g., GPUs) perform computation tasks until reaching different blocks (e.g., Multi-Head Attention or Feed-forward Neural Network), triggering new communication operations. We assume that accelerators have identical computational capabilities, processing the same amount of data at the same speed and initiating their next communication operation simultaneously.
@@
Our model also distinguishes between Data Parallelism (DP) and Model Parallelism (MP), as described in Section 2.4. We assume the initial LLM model is already stored in the accelerators’ memory, eliminating the need to exchange messages with external memory resources (e.g., NVMe disks). Instead, all communication occurs among accelerators to share information during the Forward and Backward phases in MP or to update gradients in DP. We assume five traffic-pattern configurations that may be generated in the network during the LLM training process: $C1$, $C2$, $C3$, $C4$, and $C5$. From $C1$ to $C4$, we model a massive LLM that requires MP to distribute the model across multiple accelerators. The key difference between these traffic patterns is the predominant form of MP: Tensor Parallelism (TP) or Pipeline Parallelism (PP).
@@
$C1$ assumes that TP is used extensively, i.e., most block communications follow a TP pattern (e.g., AllReduce, AllGather operations). This scenario generates significant traffic due to the need to update gradients during forward and backward propagation between devices (see Section 2.4). As a result, this traffic pattern generates $20\%$ of the communication to the inter-node network and $80\%$ to the intra-node network. Next, $C2$ and $C3$ also employ MP; however, the transferred data size between devices is smaller as these configurations rely more on PP than TP. This reduces inter-node traffic, as PP primarily uses Point-to-Point operations. Specifically, these traffic patterns generate $15\%$ and $10\%$ of the traffic to the inter-node network, while $85\%$ and $90\%$ remain within the intra-node network, respectively. $C4$ uses MP without TP, relying exclusively on PP. In this case, inter-node communication is minimized (e.g., $5\%$), as fewer updates are required between stages. Consequently, intra-node network communication is $95\%$.
@@
By contrast, $C5$ does not implement MP. We assume the LLM is not large enough to require multiple accelerators. Since it fits within a single accelerator, only DP is used to accelerate training, and no inter-node communication is needed. As a result, $100\%$ communication occurs intra-node.
3.5 Intra-node overhead modeling
As described in Section 3.3, the conversion between intra- and inter-node packets introduces extra latency and overhead, which leads to communication bottlenecks [16]. Specifically, an accelerator generates a message addressed to another accelerator placed on a remote node; that message is split into several intra-node packets that are later aggregated into one or several inter-node packets, which usually are larger in size. Thus, the intra-node packet headers will waste payload space in the inter-node packets. Moreover, when an inter-node packet arrives at a given end node, it is split again into multiple intra-node packets; the header space of the latter ones introduces an overhead in the communication, regardless of whether or not those headers were transported in the inter-node packets’ payload. In the following, we analyze the modeling of this overhead by performing a set of simulations. Moreover, we formalize this overhead through linear regression to adapt the modeling to any possible network configuration.
We have performed several simulations using the same scenarios that will be later described in Section 4. We assume an intra-node network with $8$ accelerators (Accs) per end node, where each Acc NIC generates up to $512$ Gbps of traffic. We assume a $32$-node fat-tree inter-node network topology, where each end node has a NIC connected to the intra-node and the inter-node network (see Figure 3). We assume end-node NICs can inject traffic into the inter-node network at $400$ Gbps. We analyze two different intra-node packet sizes: a commonly used intra-node packet of 148B (128B corresponding to the payload and 20B corresponding to the header), and an intra-node packet of 4KB, which is the same as the one used in the inter-node network (i.e., 64B for the header and 4032B for the payload). Note that this second packet configuration allows for measuring the network performance when there is no overhead due to intra- and inter-node packet conversion.
We use traffic configurations C4 and C5 described in Section 3.4. In C4, where each Acc generates 5% of traffic addressed to the inter-node network, each Acc will send $25.6$ Gbps to the inter-node network. So, $8$ Accs per node will generate $204.8$ Gbps in total to the inter-node network, which is less than the $512$ Gbps of inter-node links and the $400$ Gbps of end-node NICs to the inter-node network. By contrast, C5 only generates traffic to the intra-node network (i.e., no inter-node traffic) to show a scenario with no conversion overhead impact. We have simulated these traffic patterns varying the traffic generation rate from a value close to 0% of the traffic load (i.e., a very small amount of packets are generated at the Acc’s NIC) to 100% of the maximum rate that those NICs can generate (i.e., $512$ Gbps). Note that for C4, we keep 5% of intra-node packets addressed to the inter-node network regardless of the traffic load rate. Figure 4 shows the throughput results when these traffic scenarios are generated in the simulated configurations.
(a) C4 Traffic Configuration.
(b) C5 Traffic Configuration.
Figure 4: Overall Network Throughput versus Traffic Generation Load for different intra-node packets MTUs and traffic patterns in a $32$-node fat-tree inter-node network that communicates $256$ accelerators (88 accelerators per node).
[…]
Apart from the throughput metric, we have analyzed the network latency of inter-node packets and the components that compose the latency measurements. Figure 5 shows the latency results for the same scenarios described above, with packet latency divided into components (later explained in Section 4.2).
(a) Intra-node MTU of 148B.
(b) Intra-node MTU of 4096B.
Figure 5: Packet Latency divided into components versus traffic load for the traffic patterns C4 and C5 with different intra-node MTUs.
For C5, we can see that, for both MTU configurations, the latencies are negligible compared to those of C4, since there is no contention when only intra-node communication is generated. Regarding C4, we can see in Figure 5(a) that latency significantly augments when the traffic load is 70%. After saturation, several latency components consume most of the time (e.g., source intra-node, source NIC, inter-node network, and destination NIC). However, the cause of saturation happens in the destination NIC, where inter-node packets are split into intra-node packets. This process takes time and accumulates incoming traffic at destination NICs, which generates backpressure to the inter-node network that reaches source end-node NICs and their intra-node network. Note that for 148148B packets, the aggregated latency is much higher than that of 40964096B packets (see Figure 5(b)), due to the mentioned overhead 148148B packets.
On the other hand, Figure 5(b) shows that when there is no packetization overhead, the maximum latency in the communication is by far lower than in the previous case ($3.7\mu s$ versus $58\mu s$). In this scenario, note that the latency components are not affected by the Destination NIC, since there is no packetization overhead, but at the destination intra-node network, which will be the bottleneck under saturation whenever there is no overhead in the inter-to-intra-node packets conversion.
Due to the important effects of packetization overhead, we developed a mathematical analysis that establishes a relationship between the amount of inter-node traffic (i.e., the packets that accelerators send to remote end nodes) and the overhead caused by different header and payload sizes. This analysis is based on the following steps:
  1. 1.
    We selected four different mathematical functions to compare and determine if any of them fit our model (linear, quadratic, cubic, and power-law).
  2. 2.
    We collected all the necessary data sets for a 148B intra-node MTU, which includes the inter-node traffic and the throughput obtained by our simulator. We run experiments modeling a $32$-node interconnection network and eight accelerators per node, using $512$ Gbps of link speed each. So the maximum throughput per node is $4096$ Gbps.
  3. 3.
    We fit the four models to determine which one performs better and identify the possible parameters of each model. Our mathematical tool used Python packages, such as numpy and scipy, to fit the explained models. We can see the used code (i.e., modelAdjustment.py) and data from the performed simulations in the following repository [^3] https://github.com/ajtarraga/modelAdjustment/tree/main.
  4. 4.
    With this data, we evaluate the quality of each model from the four selected to determine which one represents the better option.
  5. 5.
    We select the best one and represent it.
Figure 6 illustrates this representation. After performing the mathematical analysis steps, we see that the power-law model best fits the experimental results. Based on this model, we have defined Equations 2 and 3, which allow computing the throughput behavior (considering the overhead) under different network configurations.
Figure 6: Comparison of mathematical models used to fit throughput degradation.
Equation 2 sets a relationship between the header and the payload size and between the intra- and inter-node networks. If the intra-node header impacts more than the inter-node header size, the TrafficOverhead will be higher than 1, which implies that there could be an impact (see Figure 4). However, if the impact of the intra-node header is lower than that of the inter-node ones, the value will be less than 1. If they are equal, the value will be 1, which means it does not impact.
$\text{TrafficOverhead}=\frac{\frac{\text{IntraNodeHeaderSize}}{\text{MaxIntraNodePayload}}+1}{\frac{\text{InterNodeHeaderSize}}{\text{MaxInterNodePayload}}+1}$ (2)
Equation 3 represents the function obtained when applying mathematical analysis. We have selected the power-law function as it is the one that best fits our simulation results. This formula is used to compute the expected network throughput, assuming a percentage of inter-node traffic generated at accelerators and a certain percentage of traffic overhead (computed using Equation  2).
$\text{Throughput (Gbps)}=\frac{\text{Model\_adjustment}}{\text{TrafficOverhead}\times\text{TrafficInterNode (\%)}}\times\text{NumNodes}$ (3)
Note that he Model_adjustment field has been calculated by the modelAdjustment.py script that performs the mathematical analysis applying the power-law function. The analysis of this specific value is out of scope in this paper and is left for future work.
@@ 4 Evaluation
This section describes the experiments conducted to validate our simulation model compared to a real cluster environment, and to evaluate the results of scale-up and scale-out simulations using realistic network configurations and communication patterns.
@@ 4.1 Simulation model validation
We configured our simulation model to mimic the PCIe intra-node network behavior based on the baseline end node architecture described in Section 3.1), which is available in a real computing cluster. We have used the Infiniband performance tests [17] (IB Perftests), since they generate direct communication at the user space, directly accessing the InfiniBand hardware thanks to the verbs primitives. More precisely, we measure bandwidth and latency using ib_read, ib_write, and ib_send tests with messages of sizes ranging from 128 Bytes to 4 MiB generated at the host-side user space. These messages are packetized in intra-node packets of 128B (i.e., the PCIe Maximum Payload Size or MPS in the CELLIA cluster).Table 1 shows the results for these tests run in two end nodes of the CELLIA cluster (see Section 3.1).
Table 1: Results when communicating two nodes in the real cluster
@@
Msg. Size ib_read ib_write ib_send ib_read ib_write ib_send
128 $\mathrm{B}$ 0.37 0.44 0.41 2.03 1.12 1.20
256 $\mathrm{B}$ 0.79 0.87 0.77 2.07 1.56 1.59
512 $\mathrm{B}$ 1.51 1.75 1.64 2.02 1.58 1.64
1 $\mathrm{KiB}$ 2.74 3.30 3.10 2.15 1.70 1.77
2 $\mathrm{KiB}$ 6.63 7.35 6.22 2.43 1.95 2.02
4 $\mathrm{KiB}$ 9.90 11.02 11.00 2.88 2.46 2.56
8 $\mathrm{KiB}$ 11.38 11.58 11.55 3.40 2.84 2.94
16 $\mathrm{KiB}$ 11.78 11.53 11.63 4.28 3.88 3.86
32 $\mathrm{KiB}$ 11.80 11.60 11.67 5.68 5.41 5.32
64 $\mathrm{KiB}$ 11.81 11.62 11.60 8.38 8.06 7.97
128 $\mathrm{KiB}$ 12.09 11.90 11.90 13.66 13.39 13.25
256 $\mathrm{KiB}$ 12.09 11.92 11.93 24.25 24.27 24.10
512 $\mathrm{KiB}$ 12.09 11.93 11.92 45.40 45.73 45.41
1 $\mathrm{MiB}$ 12.09 11.93 11.93 87.73 88.95 88.46
2 $\mathrm{MiB}$ 12.06 11.93 11.94 173.31 174.65 173.74
4 $\mathrm{MiB}$ 12.03 11.86 11.94 343.93 345.97 344.31
[…]
As we can see, the maximum link bandwidth for all tests is around $12.1$ GiB/s out of $12.12$ GB/s [^4] Our InfiniBand cluster uses links at EDR speed and 64b/66b encoding, so the maximum performance for 12.5GB/s is 12.12GB/s.. This maximum bandwidth is first achieved with 128128KiB messages. Note that the CELLIA cluster where we run the experiments achieves speeds of over $12.5$ $\mathrm{GiB}$$\mathrm{s}$ on the Infiniband inter-node network when transmitting raw data (i.e., without considering packet headers) and $16$ $\mathrm{GiB}$$\mathrm{s}$ on the PCIe intra-node network (see Sections 3.1 and 3.2). Similar bandwidth results can be seen in the ib_write and ib_send tests. Regarding the latency tests in the right part of the table, we can see that latency increases linearly on messages between 128128B and 128128KB. This increment is produced by the packet header overhead (i.e., 6060B) introduced when messages are split into TLPs (and later into MTUs). Note that latency skyrockets exponentially on messages larger than 128128KB because the inter-node InfiniBand link is working at full speed (i.e., 12.1 GB/s) and messages accumulate waiting to be forwarded first in the intra-node buffering space and later in the host-side memory. These results demonstrate that the inter-node network may be a bottleneck when end-node devices demand a bandwidth that is enough for the intra-node network (i.e., 16 GB/s) but not enough for the inter-node network.
@@
Figure 7: Experiment results in the CELLIA cluster and the simulator when communicating two end-node devices.
@@
To validate the intra-node network model, we modeled the ib_write Infiniband test in our simulator. We have configured the intra-node network in the simulator with the same parameters as the PCIe baseline architecture, described in Section 3.1. Figure 7 shows the bandwidth and latency results obtained running the ib_write test both in the simulator and the real cluster. Figure 7(a) compares the bandwidth results, which are virtually identical between the two series. Figure 7(b) shows the latency results that follow the same trend. These results show that, in this simple scenario, the PCIe intra-node model essentially behaves as expected. It splits the end-node CPU transactions into TLPs (i.e., intra-node packets), transferring them through the PCIe network to the NIC device. The NIC composes network packets, each with MTU=4~=4KB, which include several TLPs. These packets are sent through the InfiniBand network and, when received at the destination NIC, they are transformed again into TLPs and forwarded to the receiving buffer at the host-side user space application, where the network performance metrics are measured.